OpenStudy (anonymous):

a fielder on the gound throws a ball at an angle of 15degree to the horizontal with velocity vz at the wicket to dismiss the batsman had he thrown the ball at 45degree with a speed vb to hit wicket , then va/vb??

OpenStudy (anonymous):

So does the batsman thrown a ball at 45 degree or do we must answer it?

OpenStudy (anonymous):

I didn't get the question

OpenStudy (anonymous):

horizontal with velocity va ...sorry we have to find the ratio bet velocity

OpenStudy (experimentx):

\[ v_z = v\sin\theta \] \[ \implies v = v_z /\sin 15\] Now find the ratio of \( v_x \) and \( v_y \) at 45

OpenStudy (anonymous):

so, in each case, the ball hits the wickets is it?